In a tsunami, the height, in meters, of the highest wave, is one of the roots of the equation 2h2 - 29h - 15 = 0
a) Determine the roots of the equation 2h2 - 29h - 15 = 0 (2 marks)
b) How can you tell which of the roots will provide the answer to the problem?

Found 2 solutions by solver91311, jorel555:
Answer by solver91311(24713)



Use the quadratic formula like you would with any other quadratic. Notice that the coefficient on the constant term is negative, therefore one of the factors of the quadratic must be positive and the other negative. Since a negative number to describe the height of something is absurd, you should discard the negative root.

John

My calculator said it, I believe it, that settles it


Answer by jorel555(1290)

2h²-29h-15=0
(2h-30)(h+1/2)=0
a)h=15 or -1/2
b)Since one of the roots is negative, it can be discarded as a viable solution. ☺☺☺☺